OpenStudy (anonymous):

What is the percent of increase from 60 to 75? A. 80% B. 25% C. 20% D. 15%

OpenStudy (anonymous):

find increase then\[reqd.=\frac{ increase }{ original }*100\]

OpenStudy (jdoe0001):

from 60 to 75? well 15 units 60 + 15 = 75 so, what's 15 in percentage of 60? well, if we take say 60 as our 100% unit, and 15 as the unit of it, thus \(\large \begin{array}{cccllll} quantity&percentage\\ 60&100\\ 15&x \end{array}\bf \implies \cfrac{60}{15}=\cfrac{100}{x}\)

OpenStudy (anonymous):

Im thinking its D.

OpenStudy (anonymous):

@jdoe0001 @surjithayer

OpenStudy (jdoe0001):

well, solve for 'x", see what you get

OpenStudy (jdoe0001):

just cross-multiply and simplify

OpenStudy (anonymous):

increase=75-60=15,original=60 percent increase by above formula.